Sarah's Computer City sells 5,000 boxes of CD-RW compact discs per year. She determined that she pays $50 to process an order for discs. Her purchase price is $6 for each box of discs and she has determined that storage costs for one year are 25 percent of the purchase price. Sarah's vendor said that if she purchased 1,000 boxes at a time, she would receive a 10 percent discount. The total cost formula is TC=DP+QIP/2+DS/Q. What is Sarah's annual ordering cost if she takes advantage of the discount?
Answer #1

Total demand = 5,000 boxes

Order lot = 1,000 boxes

Number of orders = 5,000/1,000 = 5

Ordering cost per order = $50

Annual ordering cost = 5*50 = $250